OpenStudy (anonymous):

Arrange these numbers from least to greatest: 2/5, 5/7, 4/9.

OpenStudy (anonymous):

What is the LCD? :(

OpenStudy (jdoe0001):

if you had say 2 fractions say which of these 2 is bigger? 7/3 or 11/3?

OpenStudy (anonymous):

11/3

OpenStudy (jdoe0001):

right so what you can do is, grab 2 fractions at once make their denominator the same, without changing the fraction so you'd multiply both top and bottom by the same factor so let's see that \(\bf \cfrac{2}{{\color{red}{ 5}}},\cfrac{5}{{\color{blue}{ 7}}}\implies \cfrac{2}{5}\cdot \cfrac{{\color{blue}{ 7}}}{{\color{blue}{ 7}}}\quad ,\quad \cfrac{5}{7}\cdot \cfrac{{\color{red}{ 5}}}{{\color{red}{ 5}}}\) so... which one of those two is bigger?

OpenStudy (anonymous):

:/...

OpenStudy (anonymous):

Uh, that was confusing.

OpenStudy (jdoe0001):

hmmm how so?

OpenStudy (anonymous):

Why'd you change them to 7/7 and 5/5? O_o

OpenStudy (mathstudent55):

If you find the LCD of the three fractions and change each fraction to an equivalent fraction, all having the LCD as the denominator, then you can compare the numerators easily.

OpenStudy (jdoe0001):

\(\bf \cfrac{2}{{\color{red}{ 5}}},\cfrac{5}{{\color{blue}{ 7}}}\implies \cfrac{2}{5}\cdot \cfrac{{\color{blue}{ 7}}}{{\color{blue}{ 7}}}\quad ,\quad \cfrac{5}{7}\cdot \cfrac{{\color{red}{ 5}}}{{\color{red}{ 5}}}\implies \cfrac{2\cdot 7}{5\cdot 7}\quad ,\quad \cfrac{5\cdot 5}{7\cdot 5} \\ \quad \\\cfrac{14}{35}\quad ,\quad \cfrac{25}{35}\)

OpenStudy (anonymous):

@mathstudent55, I know but I can't figure out the LCD. :(

OpenStudy (the_fizicx99):

What he means to say is that you have to get both denominators to be the same, what you do to the bottom you have to do to the top so it stays equal.